# GO4Align: Group Optimization for Multi-Task Alignment

### v1

@@ 4 Related Work
Multi-Task Optimization. Multi-task optimization aims to address the task imbalance issue in multi-task learning, where each task usually has a different influence on a shared network. According to different inductive biases, we can roughly divide multi-task optimization methods into two branches: (i) Gradient-oriented methods, which solve the task balancing problem by fully utilizing the gradient information of the shared network from different tasks. Some studies report impressive performance based on Pareto optimal solutions Sener & Koltun (2018), gradient normalization Chen et al. (2018), gradient conflicting Yu et al. (2020), gradient sign Dropout Chen et al. (2020), conflict-averse gradient Liu et al. (2021), Nash bargaining solution Navon et al. (2022). However, most gradient manipulation methods usually suffer from high computational cost Kurin et al. (2022). (ii) Loss-oriented methods, which reweight task-specific losses with the help of inductive biases from the loss space, e.g., using homoscedastic uncertainty Kendall et al. (2018), task prioritization Guo et al. (2018), self-paced learning Murugesan & Carbonell (2017), similar learning paces Liu et al. (2019, 2023), random loss weight Lin et al. (2021b). Although loss-oriented methods are more computationally efficient, they often underperform gradient-oriented ones in most multi-task benchmarks. GO4Align in this work tries to trade off the overall performance and computational efficiency.
@@
Recent work Nguyen et al. (2023) weights tasks under the meta-learning setup but has lower training-time efficiency for large-scale systems with high dimensional parameter space, such as deep neural networks, limiting their applications for dense prediction tasks in MTL. The closest method to ours is the recent work FAMO Liu et al. (2023), which balances task-specific losses by decreasing task loss approximately at an equal rate. However, GO4Align proposes a new MTL optimizer that dynamically aligns learning progress across tasks by introducing group-based task interactions.
@@
Multi-Task Grouping. Multi-task grouping Standley et al. (2020); Song et al. (2022); Fifty et al. (2021) assigns tasks into different groups and trains intra-group tasks together in a shared multi-task network. Previous work Standley et al. (2020) first evaluates the transferring gains for $12^{M}-$ 1 candidate multi-task networks ($M$ is the number of tasks) and then conducts the brute-force search for the best grouping. Some works follow high-order approximation (HOA) Standley et al. (2020) to reduce the prohibitive computational cost. However, they also suffer from inaccurate estimations due to non-linear relationships between high-order gains and corresponding pairwise gains Song et al. (2022). Meanwhile, Yao et al. (2019) represents a clustered multi-task learning method, which clusters tasks into several groups by learning the representative tasks. The benefit of multi-task grouping is performance gains by training similar tasks together, and this inspires us to capture helpful group information in multi-task optimization. Also, rather than employing different multi-task networks, GO4Align introduce group-based task interactions in scaling for multi-task alignment. Moreover, we share a high-level idea of task clustering with  Thrun & O’Sullivan (1998). However, task clustering in  Thrun & O’Sullivan (1998) is limited to pairwise relationships among tasks. However, our work allows grouping-based task interactions, thus capturing more complex relationships among tasks.
5 Experiment
5.1 Multi-Task Supervised Learning
@@
Datasets and Settings. We conduct experiments on four benchmarks commonly used in multi-task optimization literature Liu et al. (2023, 2021, 2019); Navon et al. (2022). These include NYUv2 Nathan Silberman & Fergus (2012), CityScapes Cordts et al. (2016), QM9 Blum & Reymond (2009), and CelebA Liu et al. (2015). For all benchmarks, we follow the training and evaluation protocol in Navon et al. (2022); Liu et al. (2023). Detailed information about experimental results and baselines is in Appendix C.
@@
Baselines. We compare GO4Align with a single-task learning baseline, $6$ gradient-oriented methods, and $6$ loss-oriented methods. Note that single-task learning (STL) trains an independent deep network for each task. The gradient-oriented methods include MGDA Sener & Koltun (2018), PCGrad Yu et al. (2020), CAGrad Liu et al. (2021), IMTL-G Liu et al. (2020), GradDrop Chen et al. (2020), and NashMTL Navon et al. (2022). As for the loss-oriented methods, they are Linear scalarization (LS), Scale-invariant (SI), Dynamic Weight Average (DWALiu et al. (2019), Uncertainty Weighting (UWKendall et al. (2018), Random Loss Weighting (RLWLin et al. (2021a), and FAMO Liu et al. (2023).
@@
Table 1: Results on NYUv2 dataset (3 tasks). The upper and lower tables categorize baseline methods into gradient-oriented and loss-oriented types, respectively. Each experiment is repeated over 3 random seeds and the mean is reported. The best average result is marked in bold. MR and $\Delta m\%$ are the main metrics for MTL performance.
Figure 4: Efficiency comparisons on training time. Each method’s runtime is relative to a baseline method with Eq.(1), which minimizes the sum of task-specific empirical risks.
Evaluations. Following previous works (Navon et al., 2022; Maninis et al., 2019; Liu et al., 2021), we report two MTL metrics that demonstrate the overall performance over various task-specific metrics: (1) $\Delta m\%$ is the average per-task performance drop relative to STL. We assume there are $S$ metrics for all tasks. $\mathcal{M}^{s}$ denote the $s$-th metric value of a multi-task method, while $\mathcal{B}^{s}$ is the corresponding metric value of the STL baseline. Thus, we formulate the average relative performance drop as:${\bm{\Delta m}}\%=\frac{1}{S}\sum_{s=1}^{S}(-1)^{\delta^{s}}\frac{(\mathcal{M}^{s}-\mathcal{B}^{s})}{\mathcal{B}^{s}}$, where ${\delta^{s}}=1$ if higher values for the $s$-th metric are better and 00 otherwise. (2) ${\textbf{MR}}{}=\frac{1}{S}\sum_{s=1}^{S}\text{rank}(\mathcal{M}^{s})$ is the average rank of all task-specific metrics, where $\text{rank}(\mathcal{M}^{s})$ denotes the ranking of the $s$-th metric value of the model among all comparison methods. Note that in practice the lower $\Delta m\%$ and MR, the better overall performance.
@@
Table 2: Comparisons on QM9 (11 tasks), CityScapes (2 tasks) and CelebA (40 tasks). Detailed results are in Appendix D.
Effectiveness Comparisons. We provide performance comparisons on NYUv2 in Table 1. In this benchmark, our method achieves the best performance, in terms of the average performance drop, among both gradient-oriented and loss-oriented methods. Especially we find that our work is the only one that improves each task’s performance relative to the corresponding STL performance. This suggests that grouping-based task interactions can adequately alleviate the imbalance of learning progress across tasks.
@@
Efficiency Comparisons. To show the computational efficiency, in Fig. 4, we report the average training time per epoch over 5 random epochs of the proposed methods and baseline methods. We note that we run all experiments on an NVIDIA A100 and the code of baseline methods comes from  Liu et al. (2023) and  Navon et al. (2022). As shown in this figure, the proposed method GO4Align, as a new member of the loss-oriented branch, can perform more efficiently than most gradient-oriented methods. Moreover, we observe that the training time of gradient-oriented methods is proportional to the number of tasks, but our work can avoid this.
5.2 Ablation Studies
The effectiveness and efficiency of the proposed method are shown in Sec. 5.1. Next, the goal of our ablation study is to answer the following questions: (1) Can we quantify the contributions of each phrase? (2) Can we disentangle the roles of the group assignment and group weights? (3) Are there practical ways to appropriately configure hyperparameters, e.g., group number $K$?
Figure 5: Comparative analysis of the influence of group assignment and weights in the proposed AGRM principle on NYUv2. The number of groups is $2$. (a-d) have a fixed group weight $\bm{\omega}=[\omega^{1},\omega^{2}]$ but various group assignments $\mathcal{G}$ to show the influence of group assignment. (f-i) have various group weights $\bm{\omega}$ but a fixed group assignment $\mathcal{G}$ to show the influence of group weights. (e) is our method, that exploits dynamical group information from bi-level optimization. The right side of each method shows relative performance drops on each task and the average performance drop.
@@ 5.2 Ablation Studies
Contributions of Each Phase. To quantify the contributions of each phase in achieving the proposed AGRM principle on NYUv2, we report the detailed performance of our method in each phase. As shown in Table 3, compared with the scale vector in Eq. (5), the smoothness vector in Eq. (6) can compromise the performance of the “normal” and “depth” tasks, however, scarifying that of “seg.”. Based on the scale and smoothness vectors, the proposed method employs dynamical group assignment in Eq.(4) to exploit the grouping-based task interactions, thus well aligning the learning progress of similar tasks “depth” and “seg.”. We also observe that our method with both phases can improve the task-specific performance relative to STL. This demonstrates that each phase in the method complements each other, resulting in more balanced performance across tasks.
Influence of Group Assignment. To explore the influence of the group assignment matrix in the proposed AGRM principle, we assume $K=2$ on NYUv2 and make comparisons with several variants, which have various group assignments with fixed group weights. In Fig. 5, (a-c) are three fixed grouping options, e.g., (a) groups “depth" and “normal" tasks and set the group weight as $0.1$.
@@
As shown in this figure, even with fixed group weights, the appropriate group assignment shows impressive performance. In particular, grouping “Seg." and “Depth" outperforms other options. The main reason could be these two tasks are very similar to each other and far away from the “normal” task Fifty et al. (2021). We observe that variant (d) with a random grouping strategy shows lower performance than the fixed grouping options (a) and (c), which further implies the importance of appropriate group assignment in AGRM. It is worth mentioning that the proposed method in (e) without the prior information of the appropriate group assignment also captures such task correlations and each task can get performance gains compared with STL. This demonstrates that group assignment plays an important role in exploring task correlations over time in the proposed AGRM principle.
Influence of Group Weights. To study the influence of group weights in AGRM, we conduct another visualization in Fig. 5 (f-i), where we focus on various group weights with the “optimal" group assignment. (f-h) are three fixed grouping weights, e.g., (f) weight the “Normal" task in the first group with $0.6$ and weight the “Seg.” and “Depth" tasks in the second group with $0.2$. For a clear comparison, we normalize the sum of all weights to $1$.
@@
We observe that with the fixed grouping option, group weights have effects on the extent of compromising among different groups. On the NYUv2 dataset, lower weights for the first groups obtain better overall performance. The variant method (i) with random group weights achieves surprising performance, $1.75\%$, in terms of the average relative performance drop. (e) shows that our method also tends to dynamically weight the first group with a high value. This demonstrates that group weights are necessary to align the learning progress of different groups over time in the proposed AGRM principle.
@@
Figure 6: Identification of “elbow” points on NYUv2 (a) and QM9 (b). According to the conventional elbow method, we set the group number of the two datasets as 2 and 5, respectively.
@@ 6 Conclusion
Limitation. The main limitation of this work is the heuristic configuration of the group numbers. Although the search space is significantly smaller than some grouping multi-task methods Song et al. (2022), it still needs maximum $M$ runs to find the best number. Some related techniques Sinaga & Yang (2020) automatically set the group number can be added to avoid this limitation in future work.
Broader Impact. This paper is the first to consider task grouping in multi-task optimization with deep multi-task models. We propose a simple and principled way to fasten multi-task optimization with lower training-time efficiency, which has many potential societal impacts, especially in dense prediction tasks.
Appendix A Related Work on Deep Multi-task Architecture
In this section, we group multi-task learning methods into three categories: multi-task optimization, multi-task grouping, and deep multi-task architecture. The first two categories are most related to this paper and are mentioned in the main paper. To be self-contained, here we provide detailed discussions about deep multi-task architecture.
Our method GO4Align builds on multi-task grouping and multi-task optimization, inheriting advantages from both to balance different tasks in joint learning. In the following, we discuss each category of multi-task learning methods and explain how it relates to our proposed method.
Deep Multi-Task Architecture. Multi-task architecture design can be roughly categorized into either a hard-parameter sharing design Kokkinos (2017); Lu et al. (2017) or a soft-parameter sharing design Misra et al. (2016); Gao et al. (2019); Liu et al. (2019). The hard-parameter sharing design generally contains a shared encoder and several task-specific decoders. Branching points between the shared encoder and decoders are determined in an ad-hoc way Vandenhende et al. (2021, 2019), resulting in a suboptimal solution. Some work Guo et al. (2020); Lu et al. (2017); Sun et al. (2020) automatically learns where to share or branch with a network. The soft-parameter sharing design considers all parameters task-specific and instead learns feature-sharing mechanisms to handle the cross-task interactions Misra et al. (2016); Long et al. (2017). Soft-parameter sharing methods usually struggle with model scalability as the model size grows linearly with the number of tasks. GO4Align focuses on the optimization of multi-task learning, which adopts a simple hard-parameter sharing architecture as the backbone and subsequently reduces task interference in the gradient space of the shared encoder or the loss space.
Appendix B Algorithm
The pseudo-code of GO4Align is provided in Algorithm 1. For the sake of clarity, we also illustrate the optimization process in Fig. 7.
Figure 7: Optimization process of the proposed adaptive group risk minimization principle. At each iteration, given the randomly sampled mini-batch data, we first compute the group information $\bm{\omega}$ and $\bm{\mathcal{G}}$ in the lower-level optimization and then update the model’s parameter in the upper-level optimization.

### v2

@@ 4 Related Work
Multi-Task Optimization. Multi-task optimization addresses the task imbalance issue in multi-task learning, where each task usually influences a shared network differently. Task-imbalance in MTO  [1, 27] refers to imbalanced optimization rather than uneven data distributions in the task space. According to different manipulations in optimization, we roughly divide MTO methods into two branches: (i) gradient-oriented methods, which solve the task balancing problem by fully utilizing the gradient information of the shared network from different tasks. Some studies report impressive performance based on Pareto optimal solutions [18], gradient normalization [27], gradient conflicting [12], gradient sign Dropout [17], conflict-averse gradient [16], Nash bargaining solution [13]. However, most gradient manipulation methods usually suffer from high computational cost [19]. (ii) loss-oriented methods, which reweight task-specific losses with the help of inductive biases from the loss space, e.g., using homoscedastic uncertainty [5], task prioritization [28], self-paced learning [29], similar learning paces [6, 14], random loss weight [20]. Although loss-oriented methods are more computationally efficient, they often underperform gradient-oriented ones in most multi-task benchmarks. This paper tries to trade off the overall performance and computational efficiency.
@@
Recent work [30] weights tasks under the meta-learning setup but has lower training-time efficiency for large-scale systems with high dimensional parameter space, such as deep neural networks, limiting their applications for dense prediction tasks in MTL. The closest method to ours is the recent work FAMO [14], which balances task-specific losses by decreasing task loss approximately at an equal rate. However, GO4Align proposes a new MTL optimizer that dynamically aligns learning progress across tasks by introducing group-based task interactions.
@@
Multi-Task Grouping. Multi-task grouping [11, 31, 21] assigns tasks into different groups and trains intra-group tasks together in a shared multi-task network. Previous work [11] first evaluates the transferring gains for $12^{M}-$ 1 candidate multi-task networks ($M$ is the number of tasks) and then conducts the brute-force search for the best grouping. Some works follow high-order approximation (HOA) [11] to reduce the prohibitive computational cost. However, they also suffer from inaccurate estimations due to non-linear relationships between high-order gains and corresponding pairwise gains [31]. Meanwhile, Yao et al. [32] represents a clustered multi-task learning method, which clusters tasks into several groups by learning the representative tasks. The benefit of multi-task grouping is performance gains by training similar tasks together, and this inspires us to capture helpful group information in multi-task optimization. Also, rather than employing different multi-task networks, GO4Align introduce group-based task interactions in scaling for multi-task alignment. Moreover, we share a high-level idea of task clustering with  [33]. However, task clustering in  [33] is limited to pairwise relationships among tasks; meanwhile, our work allows grouping-based task interactions, thus capturing more complex relationships among tasks.
5 Experiments
5.1 Comparisons on MTL Benchmarks
@@
Datasets and Settings. We conduct experiments on four benchmarks commonly used in multi-task optimization literature [14, 16, 6, 13]: NYUv2 [34], CityScapes [35], QM9 [36], and CelebA [37]. For all benchmarks, we follow the training and evaluation protocol in [13, 14].
@@
Baselines. We compare GO4Align with a single-task learning baseline, six gradient-oriented methods, and six loss-oriented methods. Note that single-task learning (STL) trains an independent deep network for each task. The gradient-oriented methods include MGDA [18], PCGrad [12], CAGrad [16], IMTL-G [38], GradDrop [17], and NashMTL [13]. As for the loss-oriented methods, they are Linear scalarization (LS), Scale-invariant (SI), Dynamic Weight Average (DWA[6], Uncertainty Weighting (UW[5], Random Loss Weighting (RLW[39], and FAMO [14]. Detailed information about datasets and baselines is in Appendix B.
Table 1: Results on NYUv2 (3 tasks). The upper and lower tables categorize baseline methods into gradient-oriented and loss-oriented types, respectively. Each experiment is repeated over 3 random seeds, and the mean is reported. The best average result is marked in bold. MR and $\Delta m\%$ are the main metrics for overall MTL performance. Metrics with $\downarrow$ denote that the lower the better.
@@
Figure 4: Efficiency comparisons on training time. Each method’s training time is relative to a simple baseline method with Eq. (1), which minimizes the sum of task-specific empirical risks.
Evaluations. Following previous work [13, 40, 16], we report two MTL metrics that demonstrate the overall performance over various task-specific metrics: (1) $\Delta m\%$ is the average per-task performance drop relative to STL. We assume there are $S$ metrics for all tasks. $\mathcal{M}^{s}$ denote the $s$-th metric value of a multi-task method, while $\mathcal{B}^{s}$ is the corresponding metric value of the STL baseline. Thus, we formulate the average relative performance drop as:${\bm{\Delta m}}\%=\frac{1}{S}\sum_{s=1}^{S}(-1)^{\delta^{s}}\frac{(\mathcal{M}^{s}-\mathcal{B}^{s})}{\mathcal{B}^{s}}$, where ${\delta^{s}}=1$ if higher values for the $s$-th metric are better and 00 otherwise. (2) ${\textbf{MR}}{}=\frac{1}{S}\sum_{s=1}^{S}\text{rank}(\mathcal{M}^{s})$ is the average rank of all task-specific metrics, where $\text{rank}(\mathcal{M}^{s})$ denotes the ranking of the $s$-th metric value of the model among all comparison methods. Note that in practice the lower $\Delta m\%$ and MR, the better overall performance.
Table 2: Comparisons on QM9 (11 tasks), CityScapes (2 tasks) and CelebA (40 tasks). Detailed results are in Appendix C.
@@
Effectiveness Comparison. We provide performance comparisons on NYUv2 in Table 1. In this benchmark, our method achieves the best overall MTL performance among both gradient-oriented and loss-oriented methods. We observe that our work is the only one that improves each task’s performance relative to the corresponding STL performance. This suggests that grouping-based task interactions can adequately alleviate the imbalance of learning progress across tasks.
@@
Efficiency Comparison. To show the computational efficiency, in Fig. 4, we report the average training time per epoch over 5 epochs for each method. We choose LS as the relative baseline for training time (cf. RLW [39] and FAMO [14]) as it is a commonly used MTL baseline with equal weights for each task, and it does not require additional loss-oriented or gradient-oriented techniques. We note that we run all experiments on an NVIDIA A100 and the code of baseline methods comes from  Liu et al. [14] and  Navon et al. [13].
As shown in this figure, the proposed method GO4Align, as a new member of the loss-oriented branch, can perform more efficiently than most gradient-oriented methods. Moreover, when the number of tasks scales up from $2$ to $40$, the reduction in computational cost between our method and other gradient-oriented methods becomes increasingly significant, e.g., NashMTL ($2.07$) versus Ours ($1.01$) with $2$ tasks, NashMTL ($12.49$) versus Ours ($1.01$) with $40$ tasks. The main reason is that the training time of gradient-oriented methods is proportional to the number of tasks, but our work can avoid this. More experimental results are provided in Appendix C.
5.2 Ablation Study
The effectiveness and efficiency of our proposed method are shown in Sec. 5.1. Next, we answer the following questions with our ablation study: (1) Can we quantify the contributions of each phrase? (2) Can we disentangle the roles of the group assignment and group weights? (3) Can the proposed AGRM principle seamlessly integrate with existing MTO methods? (4) Are there practical ways to appropriately configure hyperparameters, e.g., group number $K$? (5) Why do we choose the K-means clustering in the proposed method?
Figure 5: Comparative analysis of the influence of the group assignment matrix and group weights on NYUv2. The x-axis in the subplots denotes the epoch, and the intensity of the color indicates the weight value. (a-d) have fixed group weights $\bm{\omega}=[\omega^{1},\omega^{2}]$ but various group assignment matrices $\mathcal{G}$. (f-i) have various group weights $\bm{\omega}$ but a fixed group assignment matrix $\mathcal{G}$. (e) is our method that dynamically exploits a group assignment matrix and group weights for each iteration. The right side of each method shows relative performance drops on each task and their average one.
@@ 5.2 Ablation Study
Contributions of Each Phase. To quantify the contributions of each phase in achieving the proposed AGRM principle on NYUv2, we report the detailed performance of our method in each phase. As the grouping is performed by Eq.(4), the first two rows in Table 3 are the variants of our method without grouping, and the last row is our method with grouping. Table 3 empirically examines the performance gains of the variant with task grouping over without grouping.
@@
In detail, compared with the scale vector in Eq. (5), the smoothness vector in Eq. (6) can compromise the performance of the “Normal” and “Depth” tasks, however, scarifying that of “Seg.”. Based on the scale and smoothness vectors, the proposed method employs dynamical group assignment in Eq. (4) to exploit the grouping-based task interactions, thus well aligning the learning progress of similar tasks “Depth” and “Seg.”. We also observe that our method with both phases can improve the task-specific performance relative to STL. This demonstrates that each phase in the method complements each other, resulting in more balanced performance across tasks.
@@
Influence of Group Assignment Matrix. To explore the influence of the group assignment matrix $\bm{\mathcal{G}}$, we assume the group number is $2$ on NYUv2 and make comparisons with several variants, which have various group assignments with fixed group weights $\bm{\omega}=[0.8,0.1]$. As shown in Fig. 5 (a-c), grouping “Seg." and “Depth" outperforms other options. The main reason could be that these two tasks are very similar and far away from the “Normal” task [21]. We observe that variant (d) with a random grouping strategy shows lower performance than the fixed grouping options (a) and (c), which further implies the importance of appropriate group assignment. It is worth mentioning that the proposed method in (e) without the prior information of the appropriate group assignment also captures such task correlations and each task can get performance gains compared with STL. This demonstrates that group assignment plays an important role in exploring task correlations over time in the proposed method.
@@
Influence of Group Weights. To study the influence of group weights, we conduct another visualization in Fig. 5 (f-i), where we focus on various group weights $\bm{\omega}$ with the “optimal" group assignment matrix $\bm{\mathcal{G}}=\{\text{Normal};\text{Seg},\text{Depth}\}$. We observe that with the fixed group assignment matrix, group weights have effects on the extent of compromising among different groups. On the NYUv2 dataset, lower weights for the first groups obtain better overall performance. The variant method (i) with random group weights achieves surprising performance, $1.75\%$, in terms of the average relative performance drop. (e) shows that our method also tends to dynamically weight the first group with a high value. This demonstrates that group weights are necessary to align the learning progress of different groups over time.
Table 4: Comparisons of existing MTO methods with the proposed AGRM on NYUv2.
MGDA 13.25 -9.11 0.83 1.38
MGDA + AGRM 6.06 -11.69 -1.14 -1.89
NashMTL -4.09 -22.01 3.15 -4.04
NashMTL + AGRM -7.75 -20.14 3.60 -4.20
FAMO -1.65 -20.02 1.29 -4.10
FAMO + AGRM 1.76 -21.17 -0.03 -4.32
GO4Align -4.03 -20.37 -1.18 -6.08
Effect of Adaptive Group Risk Minimization Principle. Empirically, the proposed adaptive group risk minimization principle (AGRM) in Sec. 3.1 can be seamlessly integrated with existing MTO methods, taking their updated task weights as group indicators. As detailed in Table 4, MTO methods combined with AGRM consistently show improved performance. MGDA with adaptive group risk minimization achieves the biggest improvement gap. Moreover, our method still outperforms others. The reason could be that the designed risk-guided group indicators are more suitable for AGRM by balancing risk scales and exploiting historical information from previous iterations.
@@
Figure 6: Identification of “elbow” points on NYUv2 and QM9. According to the conventional elbow method, we set the group number of the two datasets as 2 and 5, respectively.
@@
Effect of different clustering methods. In our main experiments, we employed standard K-means for instantiation. K-means is a widely used clustering approach. To investigate the effect of different clustering methods, we evaluate the impact of using alternative clustering algorithms.
Specifically, we tested our proposed method on NYUv2 by substituting K-means with SDP-based clustering [41] and spectral clustering [42]. As demonstrated in Table 5, these alternative clustering methods also outperform state-of-the-art approaches (FAMO, $-4.10\%$), particularly by enhancing the performance of each task over STL. Interestingly, our experiments show that the K-means clustering algorithm outperforms spectral and SDP-based clustering methods.
Table 5: Comparisons between different clustering methods in the proposed framework.
Methods $\bm{\Delta}_{seg.}\%$downarrow $\bm{\Delta}_{depth}\%$downarrow $\bm{\Delta}_{normal}\%$downarrow $\bm{\Delta}_{m}\%$downarrow Relative runtime $\downarrow$
SDP-based clustering [41] -2.97 -18.76 -1.09 -5.44 1.20$\times$
Spectral clustering [42] -1.78 -18.58 -0.06 -4.56 1.17$\times$
Ours -4.03 -20.37 -1.18 -6.08 1.02$\times$
@@ 6 Conclusion
Limitation. The main limitation of this work is the heuristic configuration of the group numbers. Although the search space is significantly smaller than some grouping multi-task methods [31], it still needs maximum $M$ runs to find the best number. Some related techniques [43] automatically set the group number can be added to avoid this limitation in future work.
Broader Impact. This paper is the first to consider task grouping in multi-task optimization with deep multi-task models. We propose a simple and principled way to fasten multi-task optimization with better training-time efficiency, which has many potential societal impacts, especially in dense prediction tasks. We provide the code for our method to encourage follow-up work. [^2] https://github.com/autumn9999/GO4Align
Acknowledgment
This work is financially supported by the Inception Institute of Artificial Intelligence, the University of Amsterdam and the allowance Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy.
Appendix
Appendix A Algorithm of GO4Align
The pseudo-code of GO4Align is provided in Algorithm 1. For clarity, we also illustrate the optimization process in Fig. 7.